OpenStudy (anonymous):

Which equation is related to this equation? w ÷ 9 = 14 A. w = 14 ÷ 9 B. w = 14 • 9 C. w = 14 – 9 D. w = 14 + 9

OpenStudy (anonymous):

oops 1 second let me change the picture

OpenStudy (anonymous):

Which equation is related to this equation? w ÷ 9 = 14 A. w = 14 ÷ 9 B. w = 14 • 9 C. w = 14 – 9 D. w = 14 + 9

OpenStudy (anonymous):

ugh

OpenStudy (anonymous):

there

OpenStudy (anonymous):

what grade are you in

OpenStudy (anonymous):

8...

OpenStudy (anonymous):

lucky.... the screen shot has the question btw

OpenStudy (anonymous):

w=5

OpenStudy (anonymous):

https://www.mathway.com/

OpenStudy (anonymous):

thats not an answer though...

OpenStudy (anonymous):

ya i know i just want you to know w+9=14

OpenStudy (anonymous):

ohh

OpenStudy (anonymous):

ya

OpenStudy (anonymous):

im still confused though on how to find the answer

OpenStudy (anonymous):

Oh wait

OpenStudy (anonymous):

Is it C?

OpenStudy (anonymous):

5=14-9? c

OpenStudy (anonymous):

CORRECT

OpenStudy (anonymous):

Yay thank you... Can i P.M you if i need more help?

OpenStudy (anonymous):

yes of course

OpenStudy (anonymous):

Thank you :D :D :D :D :D :D :D :D : D :D :D :D :D